North Eastern Locomotive Preservation Group

  • Increase font size
  • Default font size
  • Decrease font size
Q6 Blog

A record of events, activities and our work with the Q6.

As the Covid-19 pandemic had reduced the requirement for locomotives on the North Yorkshire Moors Railway, in June 2020 a small group of volunteers, working within the Covid-19 protocol, made an early start on the plan to lift the boiler out of the frames. This was to allow an external examination of the boiler by the insurance company, and to allow the boiler ticket to be extended to 2028 - 10 years from the date of the internal examination. The opportunity was taken at this time to service the axle boxes, horns and related "bottom end" components.

On 27th October 2021 the Q6 made 2 light engine test runs to Goathland wth no problems. On Thursday 28th it made another light engine trip to Goathland and then took the 14.30 train to Pickering and return with no problems. Many thanks to all those, both NELPG and NYMR, who have worked so hard over the last two years.

Over the winter we have fitted atomisers to assist with cylinder lubrication, though these have become the 2022 Q6 saga.



23rd and 24th February 2021

Ian reports two rather mild windy days with an occasional shower at Grosmont this week, with Bill Dobson, Ian Pearson, Nigel Bill, Jon Bradley and Gordon Wells on Tuesday, and Bill Dobson, Ian Pearson and Ed Bolam on Wednesday.

The following jobs were done over the two days:

  1. Right hand connecting rod fitted to the crosshead and secured.
  2. Two new cotters fitted to the crosshead little end castellated nuts.
  3. Five sections of the cab backplate undercoated.
  4. Righthand piston fitted into the cylinder using the engine lift and lifting table.
  5. Lubricator, atomisers and lefthand piston rings loaded into Jon Bradley’s car for delivery to Matthew Storey's workshop for repairs and rectification, and new fittings on the atomisers.
  6. New taper pin split, cut, and fitted to the righthand piston head nut.
  7. Left and right little end oil pots fitted (see photo).
  8. New steel angle brackets cut and drilled to be fitted to the frames to enable the oil pipes to connect to the axle boxes (see photo).
  9. One of the brackets was temporarily fitted by Bill Dobson after drilling and tapping (see photo).
  10. Lefthand piston head nut cotter cut and fitted.
  11. Righthand piston packing fitted but not yet secured.

New angle brackets for fitting to frames to enable the oil pipes to connect to the axle boxes - Ian Pearson

One of the new angle brackkets for the oil pipes temporarily fitted after drilling and tapping on 24 February 2021 - Ian Pearson

Left hand oil pot fitted on 24 February 2021 - Ian Pearson

Elsewhere, as Chris Henwood has reported separately, there was a mega shunt at the MPD yesterday to get the J27 on to the back of No 4 Road so the buffer beam can be removed and work started on the cylinder rebore (see photo). A reminder therefore that there will be a J27 working party this Saturday 27 February to progress this work. If you are willing and able to help, please contact Chris direct on 07929 733934 - he would love to hear from you! On Saturday there will also be a working party on the K1 - please contact Paul Hutchinson direct on 07964 988551 if you want to join him.

65894 standing on No 4 Road at the MPD on 24 February 2021 to allow the buffer beam to be removed and the cylinders rebored - Ian Pearson

On the Railway itself, as many of you will be aware from Chris Price's recent video message, the NYMR is planning to reopen from 1 May.

 

16th and 17th February 2021

A fine, mild, day for a change at Grosmont on Tuesday for Bill Dobson, Nigel Bill, Jon Bradley and Ian Pearson, although the first job of the day remained lighting the stove.

All the lifting equipment was inspected, including 5 strops and one shackle from the K1 support coach. The only item to fail was the pull lift from Deviation Shed. This had numerous defects so has been scrapped. Nigel B is going to acquire a new one. The current colour code is now blue.

Bill and Gordon Wells spent all morning discussing and planning what type of fittings and components are required for the new Q6 lubrication system. Last Thursday, Ian went to Middlesbrough and picked up six 18 flexible pipes from Northern Hydraulics which will be used to feed the leading, intermediate, and driving axle boxes. Two 36” flexible pipes will be ordered for the trailing axle boxes shortly. It is intended to fit atomisers to the valves. Bill has acquired two old sets of atomisers but they will need testing for leaks. These were cleaned on the outside by Nigel B using the shot blaster at the MPD. It was also found that some of the threads which will connect the oil feeds have different threads. This will make fitting a bit more complicated and an adapter will have to be made. One of the plug fittings was missing - it was lost in the workshop some three years ago and thought to be somewhere underneath the workbench .

Jon primed the five sections of cab back plate fittings, and during the afternoon the left hand little end pin was fitted and secured.

Deviation Shed new clock from Ed Bolam - 17 February 2021 - Ian Pearson

Wednesday morning started a bit dull and overcast, albeit mild, but became much brighter later. Bill wasn't feeling very well so only Ian Pearson and Ed Bolam were present. The workshop clock had stopped working early last year (not a flat battery) and today Ed brought a new one and fitted it to the south end wall (see photo). Ian told Ed about the missing plug fitting from one of the atomisers which had been lost underneath the workbench. Ed bravely cleaned all the rubbish from under the work benches and eventually found the missing plug. He then continued to clean the stubborn dirt that was left on the atomisers'

body and threads. Ian spent most of the day putting a slight chamfer on the piston heads.

All the fire extinguishers had their yearly check - usually carried out in August but the company changed its routine check to February. Bill did call in with Zak the dog later in the afternoon as he was feeling better.

 

9th and 10th February 2021

Another misleading subject heading this week as Tuesday's working party had to be cancelled due to poor weather conditions. However, in spite of the poor road conditions, Ian Pearson and Bill Dobson made it to Grosmont on Wednesday, a bitterly cold day with the odd hail and snow showers. Not surprisingly, the first job was to get the stove lit and warmed up.

The left hand connecting rod fitted on 10 February 2021 with packing pending fitting of new bearing pads - Ian Pearson

The main job of the day was fitting both connecting rods using the new hydraulic lifting table. This made the job much easier than in the past, when lots of timber packing and manpower were required. Unfortunately the bearings could not be fitted, as new felt pads had had to be made for the little ends and big end bearings, and these are presently soaking in lubricating oil. Wooden packing pieces have therefore been put around the big end crank pins to keep the connecting rod from touching the crank pins (see photo). The little end pins and big end bearings will be fitted next week.

Bill has ordered 60 metres of copper pipe for the new axle box lubrication system using the 8 feed silvertown lubricator from the upstairs store in the workshop.

 

2nd, 3rd and 4th February 2021

This subject is a bit misleading as the working party on 2 February was cancelled because of the bad weather that day. However, Bill Dobson, Ed Bolam, Gordon Wells and Ian Pearson gathered at Grosmont on a wet, cold and dismal Wednesday. The water level in the river was apparently at 4ft.

Sean Bowler informed Bill and Ian, while down at the MPD, that the crossheads were ready for fitting. Ed and Ian then took each set of crossheads from the machine shop to Deviation Shed and made a start on fitting the left side first, using our new lifting table. This makes the lifting much easier for old codgers - just jack the table up with no wooden packing required. Unfortunately though, the crosshead proved to be a bit on the tight side. As Sean had come up to Deviation Shed to see how the fitting was progressing, he and and Bill took the crosshead down and had to do some scraping of the white metal on the top slipper before refitting. This took several hours to get a good fit.

Sean Bowler working on the right hand sllde bar before fitting the crosshead - Ian Pearson

Meanwhile, Ian and Ed fitted the two remaining sand pipes to the right hand side. Bill brought the spare mechanical lubricator down from the workshop upstairs store and put it onto the right hand side running board where it is intended to be fitted (see photo). Gordon came in later to check the lubricator and mark out where some of the pipe work will go. It was planned to start fitting the connecting rods, so Ed cleaned both sets of bearings and fittings, but that was as far as it got on the day. Later on in the afternoon, Sean completed fitting the left side crosshead and made a start on the right hand side. This was also tight and more scraping of white metal and some dressing of the slide bars was required (see photo). Finishing was left for Thursday.

The mechanical lubricator on right hand running plate for marking out pipe runs - Ian Pearson

Tender and locomotive were finally moved to the rear of No 7 Road to allow enough room to get the K1 in front of the Q6 and still provide bridge access between each locomotive. The K1 and support coach were parked near the water column awaiting disposal to allow cooling down before shunting into Deviation Shed.

Today, Ian was at Grosmont again, assisting Sean Bowler with the right hand side crosshead. The crosshead is now fitted and moved along the slide bars quite easily, after dressing the slide bars and also a small amount of white metal from the top slipper. The K1 was disposed on the pit by Paul and Angie and then shunted into Deviation Shed (see photo) after the B1 was moved out, with the support coach outside on No 7 Road.

 

26th and 27th January 2021

Ian Pearson reports a fine bitter cold day at Deviation Shed on Tuesday, with Bill Dobson, Ian Pearson, Gordon Wells and Nigel Bill. Ian's first job was getting the stove lit on this very cold morning. He then assisted Nigel with checking all our lifting equipment, and recording what we own for our asset register. Nigel also liaised with Charlie Dore who looks after all the MPD’s lifting equipment, including ours, and keeps their records. The next lifting equipment inspection is due week commencing 15th February.

When Gordon arrived, he and Bill set about discussing the new lubrication plan, to feed oil to the axle boxes using an 8 feed mechanical lubricator. They then measured up for copper pipe and flexible hoses, having decided where the lubricator should be situated and where the drive to the new lubricator should come from. As it was a fine day Ian took the opportunity to dress the old paint from the left hand shoulder cladding sheet and three smaller back head cladding sheets. All are now ready for a coat of primer.

Later in the afternoon the J21 chimney was loaded onto the NYMR flat bed truck for delivery to the Great Central Railway on Wednesday for the Locomotive Conservation and Learning Trust, which is where the overhaul of their J21 (65033) is being carried out. Our thanks to the NYMR for delivering this while making a separate visit to the area.

By way of a contrast, a damp dismal day greeted Bill Dobson, Ed Bolam and Ian Pearson at Grosmont on Wednesday.

Ian discussed with Bill the need for some new taper pins for the side rod end caps and big end cap. These are going to be ordered shortly, along with a few spares. It was then on with fitting the brake blocks and sand pipes, so Ed and Ian went down to the brake block storage area at the MPD and brought 6 blocks up to Deviation Shed. All three then spent all morning fitting the 6 brake blocks - a very heavy job for old codgers I am told.

After lunch Ed and Ian started fitting the sand pipes back on the Q6. By the end of the day, all 4 pipes had been fitted to the left hand side and 2 on the right hand side. Bill completed fitting the left hand rear steam chest cylinder cock, which needed a copper washer to secure it in the correct position.

Down at the MPD, steam tests were carried out on Lambton Tank No 29 and Standard Tank No 80136.

 

14th to 20th January 2021

On Tuesday 19th, it was raining all day, but a lot milder, for the Q6 working party of Bill Dobson, Ian Pearson and Jon Bradley, who had plenty to do in Deviation Shed. Bill was busy gapping the piston rings - a long and laborious job, cutting and filing to achieve the 1/8” gap required. Ian and Jon spent the day adjusting the pipe work so that the unions connected squarely, and fitting a bracket onto the steam heat and vacuum pipes.

It was still raining on Wednesday, with Ed Bolam joining Bill Dobson and Ian Pearson. Bill first went down to the MPD to check on progress with the Q6 cross heads: after re-metalling, a start had been made on the necessary machining. Bill then spent the rest of the day completing the gapping of the piston rings. After Ed had cleaned and lit the stove, he set up a bench with Ian at the south end of Deviation shed, and dressed and primed the left hand cab shoulder cladding sheet. However, he only dressed the one piece due to the amount of dust produced while removing the old paint. Definitely an outside job for a fine day! He also cleaned and oiled the slide bars, and oiled inside the left hand cylinder. Meanwhile, Ian fitted the right hand side rod securing nuts and taper pins.

 

7th to 13th January 2021

A full week of activity to report -

7 January: Q6 - Ian Pearson and Bill Dobson met at lunchtime at Deviation Shed with the boiler in steam down at the MPD. However, Chris Kelly informed them just after lunch that he had had to fail the in house steam test due to the tea tray studs leaking. Once the boiler had cooled they would be removed. The insurance out of frames steam test arranged for the Monday was therefore cancelled and the boiler inspector informed.

12 January: Q6 - it was a cold frosty morning at Grosmont on Tuesday for Bill Dobson, Nigel Bill, Ian Pearson and Sean Bowler. Jon Bradley joined later in the afternoon. Gordon Wells also called in to rewire the kettle and water heater onto separate circuit breakers so they don’t trip out.

Bill removed the steam supply valve that feeds the vacuum brake. This was leaking steam and he found the seat required cutting. This job took most of the day, because, as is often the case, searching for cutters was a nightmare. In the end, one was borrowed from the MPD. A new brass washer is being made in place of a worn out one, so the valve will be re-fitted shortly. The water had been drained out of the boiler on Saturday morning. Nigel and Ian removed the pressure gauge and gauge glass protectors. In the afternoon, Sean completed the slide bar alignment then took both crossheads down to the MPD for measuring and checking. These will be re-metalled and machined. Later in the afternoon, with Jon assisting, the vacuum pipe was fitted.

An inspection of NELPG ladders was carried out, all of which are in Deviation Shed. One long, old, ladder has been scrapped due to defective rungs, and another has been labelled 'Not to be Used' as the integrated stabiliser has been bent and needs to be repaired. The footpath between Nos 8 and 7 road had become very cluttered with Items of the Q6 making the emergency door difficult to access. The items have been removed and the footpath is now clear.

Down at the MPD, tons of silt and ash was removed from drain sumps at the north end outside the MPD by a contract firm. We have been told that the water from Deviation Shed pit will have to be diverted away from the ash pit outside our building directly to the main drain to stop the ash getting into the traps and dumps.

13 January: Q6 - was another very cold but fine day with Bill Dobson, Ed Bolam, Ian Pearson and a late morning arrival of Nigel Bill who had picked up some consumables from Arco - 3 boxes of nitrile disposable gloves and 1 box of disposable face masks.

After getting the stove fired up, Bill, Ed and Ian fitted the steam heat pipes. They still require some securing, but are in position and also out of the way of the footpath. When Nigel arrived he spent most of the day fitting split pins to the slide bar bolts on both sides. Bill, Ed and Ian fitted fitted the front control sands rod and the left to right sands operating rod which connects to the front control rod. Ed also cleaned the the gauge glass protectors and replaced a spring on one of the protectors. Then he fitted the cylinder cocks, although some more fine adjustments are required.

There was also some tidying up in the workshop. An empty red cabinet, which has been in the way upstairs, was brought down and now sits on the bench at the south end of the workshop with a variety of tools stored in it.

Down at the MPD, a train was marshalled comprising of a well wagon with the boiler of 44806 secured to it and a brake van and Class 25 ready to move to Newbridge yard on Thursday. The boiler is going Riley engineering.

The next Q6 working parties will be on Tuesday 19 January with Ian Pearson and Bill Dobson, and Wednesday 20 January with Ian Pearson, Bill Dobson and Nigel Bill.

 

4th to 6th January 2021

This first report of the New Year covers three days. After Bill and Ian had discussed the work required to prepare for the insurance steam test on Monday 11 January, they decided to work at Grosmont on Monday this week, as well as Tuesday and Wednesday.

So, on a cold Monday morning, with heavy wintery showers, Bill and Ian met Sean Bowler at Deviation Shed to discuss what needed to be done with the tight gradient pin and bush on the leading section left hand side rod. After removing the pin which Sean ground slightly, he then dressed the inside of the bush with a small abrasive wheel. Bill and Ian started work setting up jacks under the foundation ring in preparation for lifting, to allow removal of the wooden beam and its replacement with the steel beam Bill acquired recently. But the steel beam had first to be reinforced by welding steel plate into the web of the beam, so wouldn't be ready for another 24 hours. By lunchtime Sean had finished dressing the bush. The gradient pin was tried inside the bush and was a good fit. Afterwards the front section of side rod was assembled, and the side rods fitted back onto the wheels.

Tuesday was another cold wintry day with Bill Dobson, Nigel Bill and Ian Pearson rostered, along with Sean Bowler and Dave Brown. After discussion with Sean and Dave, the main job for Sean was to prepare for lining up the slide bars, but this would mean the Q6 frames could not be moved. So it was decided that Sean would drive the tele handler and take the pallet of bagged shot blast material out of Deviation Shed to covered accommodation in a metal container near the oil store. After this, the Q6 was shunted outside and the steam reverser was taken by the tele handler to the south end of Deviation Shed and fitted back onto the Q6 frames which were then shunted back into Deviation Shed. Sean and Dave were then able to start setting up their equipment for lining up the slide bars. Unfortunately, Sean O'Driscoll's law continues to apply in 2021, and some of the fittings were found to be not compatible with the Q6: Sean had to make some new fittings. While that was being done, Dave fitted some split pins on the horn guide bolts that never got done when the frames were down at the MPD. Bill and Ian started securing the steam reverser which they planned to complete the next day. After lunch, Bill, Ian and Nigel set about removing the wooden beam from under the firebox foundation ring (see photo) using strops and a pull lift. Eventually it was replaced with a steel beam (see photo).

63395 wooden beam being removed from under the foundation ring by Nigel Bill and Bill Dobson on 5 January 2021 - Ian Pearson

63395 replacement steel beam fitted under the foundation ring on 5 January 2021 - Ian Pearson

It was no warmer on Wednesday for Bill Dobson, Ian Pearson and Ed Bolam, again, along with Sean Bowler and Dave Brown. Sean and Dave spent the day working on the left hand slide bar alignment (see photo). Bill also discussed the alignment with Sean and established that the cross heads will also have to be re-metalled. Bill has measured the boiler barrel and firebox area for insulation which will be ordered shortly, and also for new flexible axle box oil pipes, of which 16 will be required, 18” long. Ian and Ed completed fitting the steam reverser securing bolts and split pinning. In the morning, Chris Kelly fitted the bottom left hand mud hole door. This had been removed to let the boiler water out on the Tuesday before Christmas. The boiler was then filled half full and after lunch Ed and Ian set, and lit, a warming fire in the Q6. Ed also fitted both cab bottom steps which were removed when the frames were on No 4 Road in the MPD to allow the jacks to be fitted when the frames were lifted.

63395 with Sean Bowler and Dave Brown lining up the left hand slide bars on 6 January 2021 - Ian Pearson

 

22nd December 2020

Ian Pearson reports that it was a cold and frosty start for the last working party of 2020 at Grosmont yesterday, comprising Bill Dobson, Nigel Hall, Nigel Bill and Ian himself.

Nigel H got the stove going, which was very welcoming on this cold morning, before he started painting. All four eccentric rods have now had a coat of red gloss, and he also did some touching up between the frames. Nigel B completed the monthly and yearly defibrillator checks.

The main job of the day though was to move the Q6 frames from the rear of No 4 road at the MPD back into Deviation Shed. Ian spent most of the morning shunting with Charlie Wood and Chris Cubitt driving the 04 shunter D2207. By lunchtime the Q6 frames were moved into No 7 Road in Deviation Shed, with the B1 in front and 5428 Eric Treacy moved to the wheel drop shed.

Nigel B and Ian then cleaned all the old cardboard from the 4 foot area where the Q6 frames stood, which had been used on the oily floor while everyone worked between the frames, to keep them a little cleaner.

Gordon Wells, who was working down at the MPD, came in and repaired a couple of lights in Deviation Shed above the West side roller shutter door. 2 fluorescent lights were repaired, one with a new tube and the other with a new starter. Two roof lights near the stove area have been out for a while, and Gordon found that they had only tripped out. The circuit breaker was switched back in and they are now all working OK.

63395 with left hand rods removed in Deviation Shed on 22 December 2020 - Ian Pearson

After lunch, a concentrated effort by all four saw the removal of the left hand side rods (see photo) to allow the removal of the leading gradient pin which was found to be tight in the bush. The pin was eventually removed, but the bush, which was proud and also required removal, proved very stubborn, and, after trying various methods of extraction, by late afternoon it was decided to leave it until 5 January 2021 when the next Working Party is scheduled, involving Bill Dobson, Ian Pearson, Nigel Bill, and Nigel Hall, followed by Wednesday 6 January with Bill Dobson, Ian Pearson, Steve Hyman and and Ed Bolam, coronavirus and tier restrictions permitting.

In the meantime, Bill has arranged with the boiler inspector for the out of frames steam test to be held on 11 January at 1pm.

Following the retirement of Bryan Orange as Deviation Shed caretaker after 10 years in charge (he is taking on the Chairmanship of the HRA member services Training and Development Committee in the New Year), I can report that Ian Pearson and Nigel Bill have volunteered to take up the mantle of responsibility as joint caretakers for the building. My thanks to them for so doing, and also to Bryan Orange for all the hard work he has put in over the years in looking after the building and making numerous improvements to its fabric and facilities from which we are all benefitting. He will be continuing with the NELPG JVs, but I wish him good luck with his new role.

At the MPD, Repton was on a steam test and repairs to the J27's valve covers were being carried out. The No 5 Road roller shutter door was also being repaired by N.E.S.S. following a losing argument with a locomotive. Tomorrow sees the last of the Christmas Re-imagined services from Pickering, and the four Moorlander dining trains scheduled for 31 December, and 1 to 3 January are also due to run, again, subject to any further changes to the coronovirus controls which seem to be occurring on almost a daily basis. However, the planned Optimist and Rail Trail services planned between Boxing Day and 3 January have been cancelled, and whether services operate after that date will be subject to review and decision every two weeks. Regardless, the NYMR is hoping to run trains over the February half term (13 - 21 February) with a possible special visitor. So the message is to keep an eye on the NYMR website for up to date details of what is, and is not, running in early 2021.

Finally, it just remains for me to wish you all, and in particular my special correspondent Ian Pearson, all the very best for Christmas and the New Year. 2020 has been the most unimaginable and difficult one I have ever known, so my thanks to all of you, including our NYMR colleagues in the MPD in particular, for your contribution and support, particularly over the last nine months. Although 2021 looks as if it is going to start in similar vein, let's hope that the arrival of the vaccines will herald the beginning of the end of the pandemic and a return, eventually, to a more normal life. As ever though, until that time comes, please do continue to take care and, most importantly, stay safe.

 

15th and 16th December 2020

This report really starts last Thursday (10 December), which saw the Q6 re-wheeled and all eight horn stays fitted by Sean Bowler and Dave Brown, with assistance from Nigel Bill who was on a cleaning turn for the NYMR and Ian Pearson. Since then, on Friday and Monday, Sean fitted all the springs.

On a fine mild Tuesday, Bill Dobson, Nigel Bill, Jon Bradley and Ian Pearson then made a concerted effort in fitting all the brake hangers and tables, plus all the brake rods.

Wednesday though was a miserable wet mild day for Steve Hyman, Ed Bolam, Nigel Hall, Bill Dobson and Ian Pearson. Nigel Hall spent the day painting, putting another coat of black gloss on the backplate and the steam heat pipes, whilst after lunch he rubbed down the four eccentric rods and gave them a coat of red oxide.

The rest of the team made another collective effort with Sean and Dave, splitting each set of side rods into their three sections up at the wheel drop shed. They were then dragged outside so Sean could pick them up with the tele handler and take them to the MPD where they were off loaded onto a trolley and moved to the rear of No 4 Road for reassembly and fitting. The left side was fitted first, then, after lunch, the righthand side was fitted with extra help from Chris Cubitt. This was no easy job with lots of lifting, pulling, dragging, and lifting various pieces of timber packing and moving the jacks, along with pinch barring the locomotive to line up the wheels. Eventually, by late afternoon the objective for the day was achieved and the fitting of the side rods completed. In one final task for the day, the water was drained from the Q6 boiler as the boiler inspector visit for the out of frames steam test is now provisionally arranged for 11 January 2021.

On Tuesday and Wednesday, Lucie was on steam heat duty and the S15 on the Moorlander dining train. Meanwhile, following Repton failing the previous week, a warming fire was put in the J27 on Tuesday ready for a successful steam test on Wednesday and it was in charge of the Moorlander dining train today in glorious sunshine and will be again tomorrow - but in sunshine again?

 


Page 8 of 38